# Challenges for first-principles nuclear structure: 11Li and 29F

### v1

@@ Introduction.
Figure 3: Dimensionless ratio $Q/r^{2}$ for first $3/2^{-}$ level (shaded symbols) and second $3/2^{-}$ level (open symbols) of $^{\makebox[7.97226pt][r]{$\scriptstyle 11$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{Li}$}, for the protons (left) and neutrons (right), calculated with the chiral N3LO (top) and Daejeon16 (bottom) interactions. Calculated values are shown as functions of $N_{\text{max}}$}, at fixed $\hbar\omega$} (as indicated). The experimental value for $Q_{p}/r_{p}^{2}$ [70, 71] for the $3/2^{-}$ ground state is shown for comparison (solid square).
@@
Figure 4: Decompositions by $\mathrm{SU}(3)$, jointly with the number of oscillator quanta, for the first two $3/2^{-}$ levels of $^{\makebox[7.97226pt][r]{$\scriptstyle 11$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{Li}$}, calculated with the chiral N3LO (top) and Daejeon16 (bottom) interactions, for $N_{\mathrm{max}}=8$ and ${\hbar\omega}=12.5\,{\mathrm{MeV}}$. These are shown as histograms (top) and arranged by the Bohr deformation variables corresponding to the given $\mathrm{SU}(3)$ quantum numbers (bottom). Contributions from the $0{\hbar\omega}$ (shaded circles) and $2{\hbar\omega}$ (open circles) spaces are shown. Contributions from irreps which are degenerate with respect to the $\mathrm{SU}(3)$ Casimir operator (connected by dotted lines) cannot be distinguished (and, for plotting purposes, the contribution has, arbitrarily, been distributed equally between these irreps).
Now we turn to the question of the role of deformation in the intruder states, and thus (after mixing) in the $^{\makebox[7.97226pt][r]{$\scriptstyle 11$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{Li}$} halo ground state. While deformation itself is not directly observable, the electric quadrupole moment, an observable, is strongly sensitive to deformation. In Fig. 3, we show separately the proton and neutron quadrupole moments, or, more precisely, dimensionless quantities obtained by scaling the quadrupole moments to the mean-square proton and neutron radii, respectively. Such a dimensionless ratio $Q/r^{2}$ provides a measure of the quadrupole deformation, and (at least, for axially symmetric rotational nuclei) may be approximately related to the Bohr deformation variable $\beta$ [see, e.g., Eq. (13) of Ref. [72]]. From a practical viewpoint, it is also more rapidly convergent in the NCSM calculations than either the quadrupole moment or mean square radius individually [57, 72].
Take first the proton ratio $Q_{p}/r_{p}^{2}$ [Fig. 3 (left)]. Experimental measurements of the quadrupole moment are, of course, only sensitive to this proton contribution. (The experimental quadrupole moment for $^{\makebox[7.97226pt][r]{$\scriptstyle 11$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{Li}$} shows only a slight increase in magnitude relative to $^{\makebox[3.98613pt][r]{$\scriptstyle 9$}}_{\makebox[3.98613pt][r]{$\scriptstyle$}}\mathrm{Li}$} [73, 70].) Note that the calculated quadrupole moment is negative, for both the ground and excited $3/2^{-}$ states, so increasing magnitude goes downward in the plot. For the ground state, a negative sign for the quadrupole moment is as expected for a single nucleon outside a closed shell (e.g., Refs. [74, 75]), while, for the excited $3/2^{-}$, one might associate a negative quadrupole moment with prolate deformation in a rotational $K=1/2$ band (e.g., Fig. 1 of Ref. [76]). Regardless, the proton $Q/r^{2}$ is nearly the same for the normal and intruder states; this at least suggests the proton structure is not very different between the two. Our results are in close agreement with experiment [70, 71], shown as a solid square in Fig. 3, in the calculations with either the N3LO or Daejeon16 interaction. (Previous NCSM calculations obtained similar values of $Q_{p}/r_{p}^{2}$ [55, 56], including prior calculations with the Daejeon16 interaction [57].)
Now consider the calculated neutron quadrupole moment or, rather, $Q_{n}/r_{n}^{2}$, [Fig. 3 (right)]. It is very small for the normal state ($Q_{n}/r_{n}^{2}\approx-0.2$), consistent with small deviations from a spherical filled neutron shell, but much larger for the intruder state ($Q_{n}/r_{n}^{2}\approx-1.3$), as one might expect if this is a highly deformed state arising from $2{\hbar\omega}$ excitations of neutrons across the $N=8$ shell closure. If the intruder $3/2^{-}$ state is taken to be a member of a $K=1/2$ band (with inverted spin orderings due to Coriolis staggering), a plausible speculation but likely an oversimplification given the likely role of triaxiality (discussed below), then the calculated ratio corresponds to a hefty prolate intrinsic deformation of $\beta_{n}\approx 0.6$ for the neutrons. At least, this is the behavior of the observables for the N3LO calculations [Fig. 3 (b)].
However, in the Daejeon16 calculations [Fig. 3 (d)], the situation is more subtle, as we might expect in the presence of mixing. For the ground state, the neutron quadrupole moment starts small, at low $N_{\text{max}}$}, as in the N3LO calculation, but then it starts steadily climbing, from ${N_{\text{max}}}=4$ onward. For the excited $3/2^{-}$ state, the neutron quadrupole moment attains a large value at low $N_{\text{max}}$}, as in the N3LO calculation, but then turns over and starts declining in magnitude by ${N_{\text{max}}}=10$. This reflects an intruder component (with large neutron deformation) mixing into the ground state, and a normal component (with negligible neutron deformation) mixing into the excited state. As the intruder $3/2^{-}$ continues to fall in energy, and mixing with the ground state becomes more pronounced, we can only expect this trend to continue, so that the $^{\makebox[7.97226pt][r]{$\scriptstyle 11$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{Li}$} ground state structure arises from mixing with an intruder, introducing a deformed neutron halo.
To proceed beyond the simple measure of deformation provided by $Q/r^{2}$ and more properly characterize the “shape” of the state, we decompose the wave function into components in irreps of Elliott’s SU(3) [51] and interpret them approximately in terms of the standard quadrupole shape parameters $\beta$ and $\gamma$ [52]. In Fig. 4 we display the structure of the normal and intruder $3/2^{-}$ wave functions, first in terms of excitation quanta $N_{\mathrm{ex}}$, and then, within each ${N_{\text{ex}}}\hbar\omega$ space, by SU(3) irreps, and thus location on the $(\beta,\gamma)$ plane. (Specifically, these results are for the $N_{\mathrm{max}}=8$ calculation using Daejeon16 at $\hbar\omega=12.5$ MeV; the plots for higher values of $N_{\mathrm{max}}$ are not very different.) As already noted (Fig. 2), the “normal” $3/2^{-}_{1}$ state (in this calculation) is largely $0{\hbar\omega}$, while the $0{\hbar\omega}$ contribution to the “intruder” $3/2^{-}_{2}$ is less than $5\%$.
Moreover, the shapes of these states differ significantly. The normal $3/2^{-}_{1}$ state is dominated by a nearly spherical $N_{\mathrm{ex}}=0$ (1,0) irrep of SU(3) (filled circle in the $\beta$-$\gamma$ plot, where the area of the circle is proportional to the probability), while the intruder $3/2_{2}^{-}$ state has its largest contributions from highly deformed ${N_{\text{ex}}}=2$ $(5,2)$ and $(6,0)$ irreps (open circles in $\beta$-$\gamma$ plot).
Thus, despite the numerically large calculation and a priori complex nature of the microscopic wave functions, the final picture is fairly simple: the normal state follows with the naive spherical shell-model expectations, while the intruder is both mostly two-particle, two-hole and well-deformed.
Refer to captionRefer to caption
Figure 5: (Left) Relative energies for the lowest $1/2+$ and $5/2+$ levels (shaded symbols) and lowest “intruder” $1/2+$, $5/2+$, and $5/2+$ levels (open symbols), of $^{\makebox[7.97226pt][r]{$\scriptstyle 29$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{F}$}, as a function of $N_{\mathrm{max}}$, for (a) a chiral N3LO interaction and (b) the Daejeon16 interaction. Calculated values are shown as functions of $N_{\text{max}}$}, at fixed $\hbar\omega$} (as indicated). (Right) Decompositions by $N_{\text{ex}}$} for first two $5/2^{+}$ levels of $^{\makebox[7.97226pt][r]{$\scriptstyle 29$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{F}$}, in a calculation with ${N_{\text{max}}}=6$, for the Daejeon16 interaction (with ${\hbar\omega}=17.5\,{\mathrm{MeV}}$).
Figure 6: Dimensionless ratio $Q/r^{2}$ for the first $5/2^{+}$ level (shaded symbols) and second $5/2^{+}$ level (open symbols) of $^{\makebox[7.97226pt][r]{$\scriptstyle 29$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{F}$}, for the protons (left) and neutrons (right), calculated with the Daejeon16 interaction. Calculated values are shown as functions of $N_{\text{max}}$}, at fixed $\hbar\omega$} (as indicated).
Refer to captionRefer to caption
@@
Figure 7: Decompositions by $\mathrm{SU}(3)$, jointly with the number of oscillator quanta, for the first two $5/2^{+}$ levels of $^{\makebox[7.97226pt][r]{$\scriptstyle 29$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{F}$}, calculated with the Daejeon16 interaction, for $N_{\mathrm{max}}=4$ and ${\hbar\omega}=17.5\,{\mathrm{MeV}}$. These are shown as histograms (top) and arranged by the Bohr deformation variables corresponding to the given $\mathrm{SU}(3)$ quantum numbers (bottom). Contributions from the $0{\hbar\omega}$ (shaded circles) and $2{\hbar\omega}$ (open circles) spaces are shown. Contributions from irreps which are degenerate with respect to the $\mathrm{SU}(3)$ Casimir operator (connected by dotted lines) cannot be distinguished (and, for plotting purposes, the contribution has, arbitrarily, been distributed equally between these irreps).
@@
29F may be viewed as a cousin to 11Li, shifted up by a major oscillator shell. It is another halo nucleus [77] on an island of inversion [78]. Based at least on model arguments, though not direct experimental evidence, it is also thought to be deformed [79, 80]. Our maximum model space is only $N_{\mathrm{max}}=6$, with an $M$-scheme dimension of $4.7\times 10^{9}$. Nonetheless we can see trends parallel to those in 11Li.
The energy levels and the $N_{\mathrm{ex}}$ distributions are shown in Fig. 5, with the chiral N3LO interaction calculated at $\hbar\omega=18$ MeV and Daejeon16 interaction at $\hbar\omega=17.5$ MeV. Similar to 11Li, the intruder states calculated with the softer Daejeon16 interaction are lower in energy relative to the normal states than those calculated with the N3LO interaction, and the normal and intruder excited $1/2^{+}$ states (upward triangles) already approach to within $\approx 2\,{\mathrm{MeV}}$. However, we do not yet see as strong a mixing between normal and intruder states, for the model spaces we can reach in these calculations.
The $J^{\pi}=5/2^{+}$ proton and neutron quadrupole moments, scaled by the respective mean-squared radii, are shown in Fig. 6, with open symbols for the intruder states and filled symbols for normal states. We show results here only using the Daejeon16 wave functions. The pattern is, again, very similar to that of 11Li: very little difference between the normal and intruder proton quadrupole moments, suggesting the proton deformation is nearly the same, while the normal neutron quadrupole moment is nearly zero, as expected for a filled spherical shell, and large for the intruder state, suggesting strongly deformed.
The picture of a spherical normal state and a deformed intruder is backed up in Fig. 7 by the decomposition into SU(3) irreps and interpretation via the Bohr deformation parameters. The intruder decomposition and deformation is more fragmented than for 11Li; this is unsurprising considering the lore that spin-orbit splitting, which breaks SU(3), increases with mass number $A$.
@@
The main difference from 11Li is that taking 29F to higher $N_{\mathrm{max}}$, closer to convergence, is beyond current capabilities. It may not be necessary: we have strong evidence of both the importance and the numerical challenge of intruders for these island-of-inversion nuclides. Instead, using what we have learned, it may be time to turn to alternate many-body methods. Furthermore, we suggest these nuclides as excellent tests for new calculations.
@@
We have carried out large no-core shell model calculations of the island-of-inversion nuclides 11Li and 29F, using two interactions, a chiral N3LO interation and the phase-shift-equivalent interaction Daejeon16. In a naive spherical shell model picture, these nuclides should have very simple structure: spherical filled neutron shells and one proton outside filled proton shells. Instead, the ground state and other low-lying states are inverted to be primarily, or at least strongly mixed with, deformed two-particle, two-hole intruders. While the intruders descend towards the spherical normal states with increasing model space ($N_{\mathrm{max}}$), convergence is slow, even for the softer Daejeon16 interaction, and thus are a challenge to reproduce in the NCSM. Interestingly, the intruder states may be “simple” in a group-theoretical framework, that is, dominated by a few irreps.

### v2

@@ Introduction.
Figure 4: Dimensionless ratio $Q/r^{2}$ for normal (blue, shaded symbols) and intruder (red, open symbols) $3/2^{-}$ levels of $^{\makebox[7.97226pt][r]{$\scriptstyle 11$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{Li}$}, for the protons (left) and neutrons (right), calculated with the chiral N3LO (top) and Daejeon16 (bottom) interactions. Calculated values are shown as functions of $N_{\text{max}}$}, at fixed $\hbar\omega$} (as indicated). The experimental value for $Q_{p}/r_{p}^{2}$ [71, 72] for the $3/2^{-}$ ground state is shown for comparison (solid square), with the point-proton radius deduced from the charge radius as detailed in Ref. [73].
Now we turn to the role of deformation in the intruder states, and thus, ostensibly, after mixing, in the $^{\makebox[7.97226pt][r]{$\scriptstyle 11$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{Li}$} halo ground state. The electric quadrupole moment is strongly sensitive to deformation. Here we take the dimensionless ratio $Q/r^{2}$, i.e., the quadrupole moment normalized to the mean square radius (or monopole moment). This ratio is more rapidly convergent in NCSM calculations than either the quadrupole moment or mean square radius individually [57, 73] and is the more appropriate measure of deformation for an axially symmetric rotational nucleus [74, 73]. Electromagnetic measurements of the quadrupole moment (and charge radius) are, of course, are only directly sensitive to the proton density distribution. However, in Fig. 4, we examine this ratio both for the proton observables (and thus proton deformation) and for the neutron observables (and thus neutron deformation).
Take first the proton ratio $Q_{p}/r_{p}^{2}$ [Fig. 4 (left)]. (Note that the calculated quadrupole moment is negative, for both the ground and excited $3/2^{-}$ states, so increasing magnitude goes downward in the plot.) For the ground state, a negative sign for the quadrupole moment is as expected for a single nucleon outside a closed shell (e.g., Refs. [75, 76]). For the excited $3/2^{-}$ state, if it is taken to be a member of a $K=1/2$ band (with inverted spin orderings due to Coriolis staggering), such a negative (spectroscopic) quadrupole moment might be associated with prolate intrinsic deformation (e.g., Fig. 1 of Ref. [77]), at least, to the extent that it might be meaningful to associate a “deformation” to the density distribution generated by just three protons. Regardless, the proton $Q/r^{2}$ is nearly the same for the normal and intruder states ($Q_{p}/r_{p}^{2}\approx-0.6$); this at least suggests the proton structure is not very different between the two. Our results, in the calculations with either the N3LO or Daejeon16 interaction, are in close agreement with experiment [71, 72] (solid square in Fig. 4). (Previous NCSM calculations obtained similar values of $Q_{p}/r_{p}^{2}$ [55, 56], including already with the Daejeon16 interaction [57].)
The calculated neutron quadrupole moment ratio $Q_{n}/r_{n}^{2}$ [Fig. 4 (right)] is much smaller for the normal state ($Q_{n}/r_{n}^{2}\approx-0.2$), as one might expect for small deviations from a spherical filled neutron shell, but much larger for the intruder state ($Q_{n}/r_{n}^{2}\approx-1.3$), as expected for a more highly deformed state arising from excitations of neutrons across the shell closure. At least, this is the behavior of the observables for the N3LO calculations [Fig. 4 (b)].
However, in the Daejeon16 calculations [Fig. 4 (d)], the situation is more subtle, as we might expect in the presence of mixing. For the ground state, the neutron quadrupole moment starts small, at low $N_{\text{max}}$}, as in the N3LO calculation, but then it steadily climbs, from ${N_{\text{max}}}=4$ onward. For the excited $3/2^{-}$ state, the neutron quadrupole moment attains a large value at low $N_{\text{max}}$}, as in the N3LO calculation, but then turns over and starts declining in magnitude after ${N_{\text{max}}}=8$. This reflects an intruder component (with large neutron deformation) mixing into the ground state, and a normal component (with negligible neutron deformation) mixing into the excited state.
To proceed beyond the simple measure of deformation provided by $Q/r^{2}$ and more properly characterize the “shape” of the state, we decompose the wave function into components in irreps of Elliott’s $\mathrm{SU}(3)$ [51] and interpret them approximately in terms of the standard quadrupole shape parameters $\beta$ and $\gamma$ [52]. In Fig. 3 (right), we display the structure of the normal and intruder $3/2^{-}$ wave functions, decomposed first in terms of excitation quanta $N_{\mathrm{ex}}$, and then, within each ${N_{\text{ex}}}\hbar\omega$ space, by $\mathrm{SU}(3)$ irreps $(\lambda,\mu)$. These are arranged by their location on the $(\beta,\gamma)$ plane, where the area of each circle is proportional to the probability (i.e., contribution to the wave function norm) from that given $N_{\text{ex}}$} and $(\lambda,\mu)$. (Specifically, these results are for the $N_{\mathrm{max}}=8$ calculation using Daejeon16 at $\hbar\omega=12.5$ MeV; the plots for higher values of $N_{\mathrm{max}}$ are not very different.) As already noted [Fig. 3 (left)], the normal $3/2^{-}_{1}$ state (in this calculation) is largely $0{\hbar\omega}$, while the $0{\hbar\omega}$ contribution to the intruder $3/2^{-}_{2}$ is less than $5\%$.
Moreover, the shapes of these states differ significantly. The normal $3/2^{-}_{1}$ state is dominated by the nearly spherical unique $0{\hbar\omega}$ irrep $(1,0)$ of $\mathrm{SU}(3)$ (blue, shaded circle in the $\beta$-$\gamma$ plot), while the intruder $3/2_{2}^{-}$ state has its largest contributions from highly deformed $2{\hbar\omega}$ $(5,2)$ and $(6,0)$ irreps (open circles in $\beta$-$\gamma$ plot). The $(6,0)$ irrep represents a prolate, axially symmetric shape (or as close to this as can be attained in the $2{\hbar\omega}$ model space for $^{\makebox[7.97226pt][r]{$\scriptstyle 11$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{Li}$}), but the $(5,2)$ irrep lies just to the prolate side of the dotted line indicating maximal triaxiality ($\gamma=30^{\circ}$). Thus, the $\mathrm{SU}(3)$ decomposition in Fig. 3 (top right) indicates significant deviations from axial symmetry for the intruder state.
Thus, despite the numerically large calculation and a priori complex nature of the microscopic wave functions, the final picture is fairly simple: the normal state follows the naive spherical shell-model expectations, while the intruder is both mostly $2{\hbar\omega}$ (ostensibly two-particle, two-hole) and well-deformed.
Figure 5: Relative energies for the normal $1/2+$ and $5/2+$ (blue, shaded symbols) and first intruder $1/2+$ through $5/2+$ (red, open symbols) levels of $^{\makebox[7.97226pt][r]{$\scriptstyle 29$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{F}$}, for (a) a chiral N3LO interaction and (b) the Daejeon16 interaction, shown as functions of $N_{\mathrm{max}}$ (at fixed $\hbar\omega$}, as indicated).
@@
Figure 6: Decompositions by $N_{\text{ex}}$} (left) and by $\mathrm{SU}(3)$ jointly with $N_{\text{ex}}$} (right), for the normal (bottom) and intruder (top) $5/2^{+}$ levels of $^{\makebox[7.97226pt][r]{$\scriptstyle 29$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{F}$}, for the Daejeon16 interaction, in calculations with $N_{\mathrm{max}}=6$ and ${\hbar\omega}=17.5\,{\mathrm{MeV}}$. The $\mathrm{SU}(3)$ decompositions shown arranged by the Bohr deformation variables corresponding to the given $\mathrm{SU}(3)$ quantum numbers, and include contributions from the $0{\hbar\omega}$ (blue, shaded circles) and $2{\hbar\omega}$ (red, open circles) spaces. Contributions from irreps which are degenerate with respect to the $\mathrm{SU}(3)$ Casimir operator (connected by dotted lines) cannot be distinguished (and, for plotting purposes, such contributions have, arbitrarily, been distributed equally between these irreps).
@@
The nuclide 29F may be viewed as a cousin to 11Li, shifted up by a major oscillator shell. It is another halo nucleus [78] on an island of inversion [79]. Based at least on model arguments, though not direct experimental evidence, it is also thought to be deformed [80, 81]. Our maximum model space is only $N_{\mathrm{max}}=6$, with an $M$-scheme dimension of $4.7\times 10^{9}$. Nonetheless we can see trends which closely parallel those found above for $^{\makebox[7.97226pt][r]{$\scriptstyle 11$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{Li}$}.
The calculated energy levels, for the same two interactions (chiral N3LO and Daejeon16) are shown in Fig. 5, again highlighting their evolution with $N_{\text{max}}$}, at fixed $\hbar\omega$} (here, ${\hbar\omega}=18\,{\mathrm{MeV}}$ and $17.5\,{\mathrm{MeV}}$, respectively). The normal” levels (blue, shaded symbols) come lowest, as expected. These are now a $5/2^{+}$ ground state and $1/2^{+}$ excited state, corresponding to finding the valence neutron in the $0d_{5/2}$ and $1s_{1/2}$ orbitals, respectively (the $3/2^{+}$ level corresponding to the $0d_{3/2}$ orbital lies somewhat higher, at $\gtrsim 10\,{\mathrm{MeV}}$, and has been omitted for clarity). Then several more closely spaced intruder levels (red, open symbols) start at much higher energy, but swoop down with increasing $N_{\text{max}}$}. The content of the wave functions in terms of oscillator excitations, shown in Fig. 6 (left) for the $5/2^{+}$ levels, resembles that seen above for $^{\makebox[7.97226pt][r]{$\scriptstyle 11$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{Li}$} in Fig. 3 (left) — a dominant contribution at $0{\hbar\omega}$ or $2{\hbar\omega}$ respectively, for normal and intruder states, but dressed with a decaying tail of contributions at higher $N_{\text{ex}}$}.
However, once again, the intruder levels begin not quite so high in energy when calculated with the softer Daejeon16 interaction [Fig. 5(b)] — in fact, the starting energy of $\approx 13\,{\mathrm{MeV}}$, for ${N_{\text{max}}}=2$, is far below the energy of twice $\hbar\omega$} one might naively expect from the shell model. And, once again, the intruder levels approach the low-lying normal levels much sooner with the Daejon16 interaction. The intruder $1/2^{+}$ is the first to closely approach its normal counterpart (upward triangles). Indeed, already by ${N_{\text{max}}}=6$, the two approach to within $\approx 2\,{\mathrm{MeV}}$. Such strong mixing as was found for the $1/2^{-}$ states in $^{\makebox[7.97226pt][r]{$\scriptstyle 11$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{Li}$} at similar separation [Fig. 2(b)] is not yet seen here, suggesting a smaller mixing matrix element ($1\,{\mathrm{MeV}}$). Nonetheless, as the intruder energies continue to fall with $N_{\text{max}}$}, we can anticipate mixing to occur, or perhaps even an outright crossing yielding an intruder ground state.
The quadrupole moments for the normal and intruder $5/2^{+}$ levels, taken again in the dimensionless ratio $Q/r^{2}$, as a measure of deformation, are shown in Fig. 7. The pattern here is, again, markedly similar to that seen above for $^{\makebox[7.97226pt][r]{$\scriptstyle 11$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{Li}$} in Fig. 4. For the proton quadrupole moments (left), there is very little difference between the normal (blue, shaded symbols) and intruder (red, open symbols) ratios, suggesting the proton structure is nearly the same (this ratio is, again, not necessarily to be interpreted as yielding a deformation, given just a single proton outside a closed shell). Then, for the neutron quadrupole moment (right), the normal neutron quadrupole moment is again nearly zero, as expected for a filled spherical shell, and large for the intruder state, suggesting strong deformation.
The picture of a spherical normal state and a deformed intruder is supported by the decomposition into $\mathrm{SU}(3)$ irreps, shown in Fig. 6 (right), again interpreted in terms of Bohr deformation parameters. The normal state [Fig. 6 (bottom right)] is, necessarily, dominated by the unique $0{\hbar\omega}$ irrep $(2,0)$, corresponding to a single nucleon in the $sd$ shell. The decomposition of the intruder state [Fig. 6 (top right)] is more fragmented than for $^{\makebox[7.97226pt][r]{$\scriptstyle 11$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{Li}$} [Fig. 3 (top right)]. The strongest contributions are from the highly deformed $2{\hbar\omega}$ $(9,2)$ and $(10,0)$ irreps, but the probablility is spread over several of the most highly deformed $2{\hbar\omega}$ irreps. These again cover a range of triaxiality, all on the prolate side of $\gamma=30^{\circ}$. The spread over irreps is perhaps unsurprising, considering the lore that spin-orbit splitting, which breaks $\mathrm{SU}(3)$, increases with mass number $A$. Moreover, from comparing Fig. 6 with Fig. 3, it is apparent that there is simply a greater density of possible $\mathrm{SU}(3)$ irreps for the nucleus to “choose” between, within a similar region of the Bohr $\beta$-$\gamma$ deformation space.
@@
While taking $^{\makebox[7.97226pt][r]{$\scriptstyle 29$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{F}$} to higher $N_{\text{max}}$}, closer to convergence, is beyond current capabilities, we already have strong evidence that the lessons learned from $^{\makebox[7.97226pt][r]{$\scriptstyle 11$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{Li}$} apply here as well. We confirm both the importance and the numerical challenge of intruders for these island-of-inversion nuclides.
Figure 7: Dimensionless ratio $Q/r^{2}$ for the normal (blue, shaded symbols) and intruder (red, open symbols) $5/2^{+}$ levels of $^{\makebox[7.97226pt][r]{$\scriptstyle 29$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{F}$}, for the protons (left) and neutrons (right), calculated with the Daejeon16 interaction. Calculated values are shown as functions of $N_{\text{max}}$} (at fixed $\hbar\omega$}, as indicated).
@@
We have carried out large no-core shell model calculations of the island-of-inversion nuclides 11Li and 29F, using two interactions, a chiral N3LO interation and the phase-shift-equivalent interaction Daejeon16. In a naive spherical shell model picture, these nuclides should have very simple structure: spherical filled neutron shells and one proton outside filled proton shells. Instead, the ground state and other low-lying states are inverted to be primarily, or at least strongly mixed with, deformed two-particle, two-hole intruders. In ab initio NCSM calculations, while the intruders descend towards the spherical normal states with increasing model space ($N_{\text{max}}$}), convergence is slow, even for the softer Daejeon16 interaction, and thus the intruders are a challenge to reproduce in the NCSM. Interestingly, the intruder states may be “simple” in a group-theoretical framework, that is, dominated by a few irreps, representing the largest possible deformations with some triaxial tendencies).